A ray of light traveling in a block of glass (n=1.52) is incident on the top surface at an angle of 57.4? with respect to the normal in the glass. If a layer of oil is placed on the top surface of the glass, the ray is totally reflected.

What is the maximum possible index of refraction of the oil?
Answer #1

apply snells law na \(\sin A=n b \sin B\)

where \(A\) is angle of incidence, \(B\) is angle of Refraction

na and \(\mathrm{nb}\) are refractive index

for criticla, angle \(\mathrm{B}=90\) deg

\(\mathrm{nb}=1.52 * \sin 57.4\)

\(\mathrm{nb}=1.28\) is the rerfractive index of oil layer